# Ultraviolet Behavior of the Wheeler–DeWitt Equation in Hořava–Lifshitz Gravity

### v1

@@ 4 Spherical and Hyperbolic Black Hole Interiors
where $d\Omega^{2}_{2}$ is the line element of the 22 dimensional sphere with unit radius, and $r_{s}$ is a constant of dimension length, identified with the horizon radius in the spherically symmetric classical solution.
@@
$\Pi_{X}$ $=(\lambda-3)\dot{X}-2(\lambda-1)\dot{Y},$ (4.4)
$\Pi_{Y}$ $=-2(\lambda-1)\dot{X}+2(2\lambda-1)\dot{Y},$ (4.5)
@@
$\Psi(X,Y)=\left\{\begin{array}[]{ll}e^{-ikY}e^{-\frac{\lambda-1}{2\lambda-1}ikX}K_{\frac{\sqrt{3\lambda-1}}{\sqrt{2}(2\lambda-1)}ik}(\gamma e^{X})&({\rm original\ action}),\\ e^{-ikY}e^{-\frac{\lambda-1}{2\lambda-1}ikX}J_{\frac{\sqrt{3\lambda-1}}{\sqrt{2}(2\lambda-1)}ik}(\gamma e^{X})&({\rm analytically\ continued\ action}),\end{array}\right$. (4.10)
@@
Although $\Psi(X,Y)$ is obtained by solving the equation in the UV limit, it is useful to compare it with the corresponding solution in general relativity. In the special case $\lambda=1$, the wave function reduces to a familiar form.
$\Psi(X,Y)=e^{ikY}K_{ik}(\gamma e^{X}),$ (4.12)
@@ 5 Planar Black hole interiors
where $d\bar{\Omega}^{2}_{2}$ is the line element of the flat 22 dimensional torus.
@@
$\Psi(Z,W)=\left\{\begin{array}[]{ll}e^{-ikZ}e^{-\frac{10(\lambda-1)}{8\lambda-11}ikW}K_{\frac{\sqrt{9(3\lambda-1)}}{\sqrt{2}(11-8\lambda)}ik}(\nu e^{W})&({\rm original\ action}),\\ e^{-ikZ}e^{-\frac{10(\lambda-1)}{8\lambda-11}ikW}J_{\frac{\sqrt{9(3\lambda-1)}}{\sqrt{2}|11-8\lambda|}ik}(\nu e^{W})&({\rm analytically\ continued\ action}),\end{array}\right$. (5.8)
@@
where AA is the normalization constant and $\sigma$ is the standard deviation of the pulse at $Z=W$. Then, the wave function is
@@ 6 Non cosmological constant Black hole interiors
$\Psi(Z,W)=\left\{\begin{array}[]{ll}e^{-ikW}e^{-\frac{20(\lambda-1)}{25\lambda-19}ikZ}K_{\frac{\sqrt{-2(400\lambda^{2}-827\lambda+409)}}{(25\lambda-19)}ik}(\rho e^{Z})&(c_{8}<0),\\ e^{-ikW}e^{-\frac{20(\lambda-1)}{25\lambda-19}ikZ}J_{\frac{\sqrt{-2(400\lambda^{2}-827\lambda+409)}}{(25\lambda-19)}ik}(\rho e^{Z})&(c_{8}>0),\end{array}\right$. (6.9)

### v2

@@ 4 Spherical and Hyperbolic Black Hole Interiors
where $d\Omega^{2}_{2}$ is the line element of the $2$ dimensional sphere with unit radius, and $r_{s}$ is a constant of dimension length, identified with the horizon radius in the spherically symmetric classical solution.
@@
$\Pi_{X}$ $=(3-\lambda)\dot{X}-2(1-\lambda)\dot{Y},$ (4.4)
$\Pi_{Y}$ $=-2(1-\lambda)\dot{X}+2(1-2\lambda)\dot{Y},$ (4.5)
@@
$\psi_{k}(X,Y)=\left\{\begin{array}[]{ll}e^{-ikY}e^{-\frac{\lambda-1}{2\lambda-1}ikX}K_{\frac{\sqrt{3\lambda-1}}{\sqrt{2}(2\lambda-1)}ik}(\gamma e^{X})&({\rm original\ action}),\\ e^{-ikY}e^{-\frac{\lambda-1}{2\lambda-1}ikX}J_{\frac{\sqrt{3\lambda-1}}{\sqrt{2}(2\lambda-1)}ik}(\gamma e^{X})&({\rm analytically\ continued\ action}),\end{array}\right$. (4.10)
@@
Although $\psi_{k}(X,Y)$ is obtained by solving the equation in the UV limit, it is useful to compare it with the corresponding solution in general relativity. In the special case $\lambda=1$, the wave function reduces to a familiar form.
$\psi_{k}(X,Y)=e^{ikY}K_{ik}(\gamma e^{X}),$ (4.12)
@@ 5 Planar Black hole interiors
where $d\bar{\Omega}^{2}_{2}$ is the line element of the flat $2$ dimensional torus.
@@
$\psi_{k}(Z,W)=\left\{\begin{array}[]{ll}e^{-ikZ}e^{-\frac{10(\lambda-1)}{8\lambda-11}ikW}K_{\frac{\sqrt{9(3\lambda-1)}}{\sqrt{2}(11-8\lambda)}ik}(\nu e^{W})&({\rm original\ action}),\\ e^{-ikZ}e^{-\frac{10(\lambda-1)}{8\lambda-11}ikW}J_{\frac{\sqrt{9(3\lambda-1)}}{\sqrt{2}|11-8\lambda|}ik}(\nu e^{W})&({\rm analytically\ continued\ action}),\end{array}\right$. (5.8)
@@
where $A$ is the normalization constant and $\sigma$ is the standard deviation of the pulse at $Z=W$. Then, the wave function is
@@ 6 Non cosmological constant Black hole interiors
$\psi_{k}(Z,W)=\left\{\begin{array}[]{ll}e^{-ikW}e^{-\frac{20(\lambda-1)}{25\lambda-19}ikZ}K_{\frac{\sqrt{-2(400\lambda^{2}-827\lambda+409)}}{(25\lambda-19)}ik}(\rho e^{Z})&(c_{8}<0),\\ e^{-ikW}e^{-\frac{20(\lambda-1)}{25\lambda-19}ikZ}J_{\frac{\sqrt{-2(400\lambda^{2}-827\lambda+409)}}{(25\lambda-19)}ik}(\rho e^{Z})&(c_{8}>0),\end{array}\right$. (6.9)